OpenStudy (anonymous):

can some one check this How many solutions does the equation have? a+5=1/5(5a+25) i think its no solution but can someone tell me if im right

OpenStudy (anonymous):

ya me too i think its no soluton

ganeshie8 (ganeshie8):

a+5=1/5(5a+25) distribute 1/5 on right hand side, what do u get ?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

how did u get 30a ?

ganeshie8 (ganeshie8):

a+5=1/5(5a+25) distribute 1/5 on right hand side, you get : a+ 5 = 1/5*5a + 1/5 * 25

ganeshie8 (ganeshie8):

simplifying, you get : a + 5 = a + 5

OpenStudy (anonymous):

oh thanks

ganeshie8 (ganeshie8):

which is true for all "a"

ganeshie8 (ganeshie8):

so there will be infinitely many solutions

OpenStudy (anonymous):

oh ok thanks can you help me with 3 more?

ganeshie8 (ganeshie8):

I'll try, ask..

OpenStudy (anonymous):

How many solutions does the equation have? 4x+2(x-3)=8x+12

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

what do u think ?

ganeshie8 (ganeshie8):

i mean, do u know how to solve these without guessing ?

ganeshie8 (ganeshie8):

solving these is easy

ganeshie8 (ganeshie8):

you just need to isolate x

ganeshie8 (ganeshie8):

4x+2(x-3)=8x+12 first distribute 2 for the parenthesis on left side

ganeshie8 (ganeshie8):

what do u get ?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

on this one i seriously dont even know to take a guess its hard for me

ganeshie8 (ganeshie8):

@igivemedalsandfollow!

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

its your chance to understand, it wont take any time u wil understand them in less than 5 minutes

OpenStudy (anonymous):

ok lets do it

ganeshie8 (ganeshie8):

you need to know distributive law : ` a(x+y) = ax + ay `

ganeshie8 (ganeshie8):

heard of that law before ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

4x+2(x-3)=8x+12 if you just want answer, the answer is `only one solution`. but with that you and me wont learn anything

ganeshie8 (ganeshie8):

4x + 2(x-3) = 8x + 2 distributing 2 you get : 4x + 2x - 6 = 8x + 2 combining like terms on left hand side you get : 6x - 6 = 8x + 2

OpenStudy (anonymous):

ok explain

ganeshie8 (ganeshie8):

by this time, it should be clear that there is an unique solution. if not, go ahead and solve x

OpenStudy (anonymous):

ok so here i tried doing this one and i got these 2 anwers do im almost shure i got them rightWhich expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = Choose exactly two answers that are correct. A. 14x – 15 B. 2(7x – 10) C. 15x – 14 D. 2(7x – 9)

OpenStudy (anonymous):

i choosed a and d

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

a is correct, but d is not correct

ganeshie8 (ganeshie8):

D gives you infinitely many solutions

OpenStudy (anonymous):

so c right

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

no, C gives you an unique solution

OpenStudy (anonymous):

so b ??????????????????? how

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

A and B